0.05x+0.25yh=11; 0.15x+0.05y=12 Solve by the elimination method

Answer by richwmiller(17219)


I am going to assume the h is a typo.
0.05x+0.25y=11
0.15x+0.05y=12
multiply first by 3
.15x+.75y=33
.15x+.05y=12
subtract second from first
.7y=21
y=210/7=30
.15x+.05(30)=12
.15x+1.50=12
15x+150=1200
15x=1050
x=1050/15
x=70